OpenStudy (anonymous):

How would I write this equation in scientific notation? (2.6 × 10–4)(2.2 × 108)

OpenStudy (anonymous):

Is 5.72 x 10^-32 Right?

OpenStudy (anonymous):

@Phil1111

zepdrix (zepdrix):

What is that last number supposed to be? Is that 10^8?

OpenStudy (anonymous):

Yeah and the other one is 10^-4

zepdrix (zepdrix):

Exponent rules tell us to ADD the powers,\[\Large\rm 10^{-4}\times 10^8=10^{-4+8}\]

zepdrix (zepdrix):

The 5.72 looks correct though :)

OpenStudy (anonymous):

I multiplied the first numbers of each part

OpenStudy (anonymous):

Is that how you are supposed to do it?

zepdrix (zepdrix):

\[\Large\rm (2.6\times10^{-4})(2.2\times10^8)\]\[\Large\rm =2.6\cdot2.2\times 10^{-4}\cdot10^8\]\[\Large\rm =5.72\times10^{-4+8}\]\[\Large\rm =5.72\times10^{4}\]Yes :) Group the tens for multiplication, and group the non-tens for multiplication as I showed above.

OpenStudy (anonymous):

Thank you so much

zepdrix (zepdrix):

np c: